For each day, independent of the others, the length of time for one individual to be served at

a cafeteria is a random variable having an exponential distribution with a mean of 4 minutes. What is the

probability that a person is served in less than 3 minutes on at least 4 of the next 6 days?


1
Expert's answer
2022-05-18T14:18:46-0400

First, we find the probability that a person is served in less than 3 minutes on a single day, using the exponential distribution.

Exponential distribution:

The exponential probability distribution, with mean m, is described by the following equation:  

"f(x)=\mu e^{-\mu x}"

in which "\mu=1/m" is the decay parameter.

The probability that x is lower or equal to a is given by:  

"P(X\le x)=1-e^{-\mu x}"

In this problem, mean of 4 minutes, hence "m=4, \mu=0.25"

The probability that a person is served in less than 3 minutes on a single day is:

"P(X\le 3)=1-e^{-0.25(3)}=0.5276"

Now, for the 6 days, we use the binomial distribution:

"P(X=x)=C_{n,x}p^x(1-p)^{n-x}"

"C_{n,x}=\frac {n!} {x!(n-x)!}"

The parameters are:

  • x is the number of successes.
  • n is the number of trials
  • p is the probability of a success on a single trial.

In this problem "n=6" , "p=0.5276" .

Hence, "P(X\ge4)=P(X=4)+P(X=5)+P(X=6)"

"P(X=4)=C_{6,4}*(0.5276)^4*(0.4724)^2=0.2594"

"P(X=5)=C_{6,5}*(0.5276)^5*(0.4724)^1=0.1159"

"P(X=6)=C_{6,6}*(0.5276)^6*(0.4724)^0=0.0216"

Then, "P(X\ge 4)=0.2594+0.1159+0.0216=0.3969\approx39.7" % that a person is served in less than 3 minutes on at least 4 of the next 6 days.